# Enhanced Computationally Efficient Long LoRA Inspired Perceiver Architectures for Auto-Regressive Language Modeling

### v1

@@ 7 Conclusion
Efficient computation of attention in transformer models is an important area of research with significant impact on the training budgets of large language models. In this work, we enhance the PerceiverAR architecture by examining its limitations in language modeling and providing different enhancements to improve its performance. One of the reasons, we focus on the PerceiverAR design is that it divides the input context into two components i.e., latent and history. This two level breakdown can be exploited further in a Long LoRA style efficient computation of attention without loss of performance with respect to full attention. Similar to Long LoRA, where the different attention heads divide the input tokens into groups, and half the groups are shifted to propagate the attention, we equivalently divide the input into segments, and compute the PerceiverAR operation on a pair of consecutive overlapping segments. This overlap of half segments accomplishes the same goal as Long LoRA and causes the propagation of attention information from all previous contexts to the current segment. Our results indicate efficient computation with excellent perplexity. Finally, it should be noted that our Long LoRA inspired PerceiverAR based architecture can use the Long LoRA in the attention heads to further improve the language model performance.

### v2

@@ 7 Conclusion
Efficient computation of attention in Transformer models is an important area of research, with significant impact on the design of LLMs. We start with the PerceiverAR architecture as a basis for developing enhanced architectures. PerceiverAR divides the input into two components i.e., the history and the latent. This two level breakdown can be exploited in different ways. We first propose three distinct enhanced PerceiverAR architectural paradigms, each with their own efficiency performance trade-offs. Based on the two main limitations of the PerceiverAR (limited latent training and lossy history), and the strengths of different enhanced designs, we develop the Efficient Context Propagating Perceiver (ECP). The ECP architecture utilizes both the context and the latent sequences in autoregressive training. In addition, the ECP design operates with similar attention complexity as LongLoRA, making it extremely computationally efficient. It also implicitly extracts better information than existing architectures, via its overlapping pairwise segment attention.